OpenStudy (anonymous):

just need someone to confirm my answers....use substitution e^(x)=u to valuate the integral e^(x)/(1+e^(x)) dx

OpenStudy (anonymous):

So du=e^x dx \[\int\limits \frac{du}{1+u}\] Let s=1+u ds=du \[\int\limits \frac{ds}{s}=\ln|s|+C \rightarrow \ln|u+1|+C \rightarrow \ln|e^x+1|+C\]

myininaya (myininaya):

\[\int\limits_{}^{}\frac{du}{u}, with u=e^x+1 => du=e^x dx\]

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

\[withu\]

myininaya (myininaya):

i wonder why they want you to use u=e^x not u=e^x+1

OpenStudy (anonymous):

It said let u=e^x thats why I did it that way.

myininaya (myininaya):

i know male

OpenStudy (anonymous):

cant make a difference can it? since 1 is a number i mean

myininaya (myininaya):

i just don't know why they would ask you to use that substitution

OpenStudy (anonymous):

ya that's the problem says it's supposed to be. this teacher likes to give hard to impossible problems tho so i figured i'd double check

myininaya (myininaya):

i mean its okay but i like mine more

OpenStudy (anonymous):

you better hurry before saifoo becomes a sensei

myininaya (myininaya):

hey satellite what number of medals do i need to get to sensei level?